(a) A body of mass 60 kg rests on the floor of a lift (take g=10ms-2). Calculate:

 

(i)   the reaction force exerted on the body by the floor of the lift, when the lift is travelling at a steady speed of 5 ms-1 upwards;     


(i)   the reactive force exerted on the body by the floor of the lift, as its upward motion is brought to rest by a constant deceleration of 2 ms-2.                                               


1
Expert's answer
2021-02-23T10:03:52-0500

1)


"N=mg=(60)(9.8)=588\ N"

2)


"N=m(g-a)=(60)(9.8-2)=468\ N"